For a Saturday matinee, adult tickets cost $8.50 and kids under 12 pay only $6.00. If 80 tickets are sold for a total of $555, how many of the tickets were adult tickets and how many were sold to kids under 12?
Answer by mananth(16949)

1.00 x + 1.00 y = 80.00 .............1

8.50 x + 6.00 y = 555.00 .............2
Eliminate y
multiply (1)by -6.00
Multiply (2) by 1.00
-6.00 x -6.00 y = -480.00
8.50 x + 6.00 y = 555.00
Add the two equations
2.50 x = 75.00
/ 2.50
x = 30.00
plug value of x in (1)
1.00 x + 1.00 y = 80.00
30.00 + y = 80.00
y = 80.00 -30.00
y = 50.00
y = 50.00
x= 30 number of adult tickets
y= 50 number of child tickets